OpenStudy (anonymous):

Find the area of triangle ABC with the given parts. Round to the nearest whole number. A = 26.4° b = 12.3 in. c = 7.7 in. a- 19.1 in.2 b- 42.4 in.2 c- 44.4 in.2 d- 21.1 in.2

OpenStudy (anonymous):

in.2 = in^2

OpenStudy (alexwee123):

you want me to show work or just solve?

OpenStudy (anonymous):

Solve if you dont mind

OpenStudy (anonymous):

Im gonna end up learning it. I just wanted to knock this quiz out real quick.

OpenStudy (alexwee123):

k give me a sec

OpenStudy (anonymous):

take your time:) the quiz is 10 questions if u wanna help on all haha

OpenStudy (anonymous):

but for some reason I get 2 attempts. and after the first attempt I get to see the correct answers...

OpenStudy (anonymous):

so I really just need liike 6/7 correct so make it look like I tried. Then I get a 100 no matter what. If you get what im saying.

OpenStudy (anonymous):

area = \[\frac{1}{2} \sin \theta ab\] 21 inch is the answer

OpenStudy (anonymous):

oh dang. I didnt know that was the formula. Would figuring the sides work? Then doing 1/2 * a * b?

OpenStudy (anonymous):

Or do u have to use the formula u said?

OpenStudy (alexwee123):

21.1

OpenStudy (anonymous):